OpenStudy (anonymous):

Can someone help me on this? Can't figure out how to do this, and my grades depend on it! -10 > -2(2x - 9)

OpenStudy (danjs):

could you solve that if it were an = sign instead?

OpenStudy (danjs):

The only thing to remember here, is iff you multiply or divide by a negative, the signs on every term will switch, and you also have to switch the > < around too

OpenStudy (danjs):

if you want to multiply both sides by -1 to maybe clean it up a bit, you also have to change the sign -1*-10 < -1*-2(2x - 9) 10 < 2*(2x-9) that is the only thing to watch out for

OpenStudy (danjs):

ie -16 < 16 same thing with the -1 on that 16 > -16 see the sign has to flip to keep true

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

Yeah. I'm still trying to figure it out according to what Dan is saying.

jhonyy9 (jhonyy9):

so my opinion that you can dividing both sides by 2 - yes - so what will get ?

jhonyy9 (jhonyy9):

do you wan cooperate please ?

jhonyy9 (jhonyy9):

so you get -5> - (2x-9) yes ?

OpenStudy (anonymous):

yes.

jhonyy9 (jhonyy9):

-(2x-9) = ?

jhonyy9 (jhonyy9):

= -2x+9 = 9-2x --- yes ?

jhonyy9 (jhonyy9):

so -5 > 9-2x - subtract from both sides -9 - so what will get ?

jhonyy9 (jhonyy9):

not minus 9 just 9 - sorry

OpenStudy (anonymous):

-14 > 2x?? Sorry, really bad at this.

jhonyy9 (jhonyy9):

why +2x ?

jhonyy9 (jhonyy9):

above was 9-2x - yes ?

jhonyy9 (jhonyy9):

so what will get ?

jhonyy9 (jhonyy9):

-14>-2x - yes ?

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

divide both sides by 2 and will get -7>-x

jhonyy9 (jhonyy9):

so what have said above Dan to you ,than -7 > -x so than x > 7 hope you understand it

OpenStudy (danjs):

sorry you never responded, thought you left

OpenStudy (anonymous):

^Dan JavaScript Eh? :P

OpenStudy (anonymous):

\[-10>-4x+18\] I distributed the 2 \[-28>-4x\] I subtracted 18 on each side\[7<x\] I divided the negative 4 on each side to get the variable alone. I switched the sign because when you divide a negative on each side, you have to flip the sign. I took pre-algrebra two years ago and I hope this helps, it seems that you got the answer, but I thought this might help explain the steps.